# Finite Distance Corrections to Vacuum Birefringence in Strong Gravitational and Electromagnetic Fields

### v1

@@ Abstract
We study polarization dependent photon propagation in static, spherically symmetric spacetimes permeated by strong magnetic fields, with the aim of quantifying how finite emission and detection radii modify vacuum birefringence signals. Working in the geometric optics limit of nonlinear electrodynamics, we formulate the two polarization modes as null geodesics of distinct effective (optical) metrics. We then develop a controlled weak-coupling expansion that cleanly separates the standard gravitational deflection from the birefringent contribution induced by the electromagnetic nonlinearity. Using a finite distance Gauss-Bonnet construction on the associated optical manifolds, we derive a general expression for the differential bending angle in which the source and observer are kept at arbitrary radii, thereby extending the usual scattering-at-infinity treatment. As benchmark applications, we specialize our results to the Euler-Heisenberg effective action of QED and to Born-Infeld electrodynamics. We find that the observable birefringence is generically reduced by finite-distance truncation of the curvature flux, and we provide explicit correction series suitable for data analysis. For magnetar-motivated dipolar fields, this geometric effect yields a suppression factor that can diminish the predicted polarization-dependent deflection by as much as $\sim 50\%$ for limb emission relative to asymptotic scattering estimates. Our results furnish a necessary finite-distance calibration for interpreting current and future X-ray polarimetry measurements and for placing unbiased constraints on strong-field QED and broader NLED parameters.
@@ I Introduction
The interaction of light with strong gravitational and electromagnetic fields provides a uniquely clean arena in which fundamental physics becomes observationally testable. In classical general relativity, light propagation is achromatic and polarization blind: null rays follow the spacetime light cone, and gravitational lensing is a purely geometric effect [Einstein_2005, Dyson:1920cwa]. Quantum field theory in external backgrounds, however, predicts that the vacuum itself acquires nontrivial optical properties in sufficiently intense electromagnetic fields. In particular, vacuum polarization in QED leads to the celebrated Euler-Heisenberg effective action and to birefringence and dispersion of photon modes in strong magnetic fields [Heisenberg:1936nmg, Bialynicka-Birula:1970nlh, Adler:1971wn, Schwinger:1951nm]. Beyond QED, nonlinear electrodynamics (NLED) also arises as an effective description in broader contexts, including string-inspired constructions, and admits a rich phenomenology when coupled to gravity [Fradkin:1985qd, Tseytlin:1995uq, Sorokin:2021tge, Kruglov:2015yua].
@@
The advent of horizon-scale imaging by the Event Horizon Telescope (EHT), exemplified by the image of Sagittarius A*, has opened a new precision window for testing gravity and fundamental physics in the strong-field regime, where deviations from the Kerr paradigm can be constrained directly at the scale of the photon ring [Vagnozzi:2022moj]. In this context, nonlinear electrodynamics provides a particularly well-motivated arena: magnetically charged black holes sourced by NLED can yield shadow and lensing signatures that are, in principle, accessible to EHT analyses and can be used to bound the underlying electromagnetic nonlinearity [Allahyari:2019jqz]. More broadly, combining electromagnetic probes (shadow, lensing) with dynamical and particle-physics observables (quasinormal modes, greybody bounds, and propagation effects) offers a multi-messenger strategy to break degeneracies among model parameters [Pantig:2022gih, Lambiase:2005gt]. Recent studies of rotating Einstein-Euler-Heisenberg black holes further demonstrate how QED-inspired nonlinearities can imprint simultaneously on shadows and ringdown spectra, reinforcing the relevance of strong-field birefringent and dispersive effects for upcoming high-resolution observations [Lambiase:2024lvo]. NLED provides a well-motivated extension of Maxwell theory that captures high-field corrections arising both as effective quantum phenomena, such as the Euler-Heisenberg vacuum polarization effects underlying photon dispersion and photon splitting in strong backgrounds [Bialynicka-Birula:1970nlh, Adler:1971wn], and as fundamental structures emerging in broader frameworks, including string-inspired constructions [Fradkin:1985qd, Tseytlin:1995uq]. When coupled to general relativity, NLED generically alters the near-horizon and strong-curvature regimes and has long been known to admit black-hole geometries whose properties can deviate markedly from the Reissner-Nordström solution [Pellicer:1969cf, deOliveira:1994in, Yajima:2000kw, Ruffini:2013hia, Kruglov:2015yua]. A particularly influential development is that suitable NLED sources can yield regular (nonsingular) black holes, providing controlled arenas to explore the interplay between energy conditions, effective stress tensors, and astrophysical observables [Ayon-Beato:1998hmi, Bronnikov:2000vy, Dymnikova:2004zc, Balart:2014cga, Fan:2016hvf, Bambi:2013ufa, Li:2024rbw]. These geometries have renewed relevance because NLED-induced modifications can imprint themselves on wave and ray probes (e.g., quasinormal spectra, birefringence, and optical signatures such as weak deflection and shadow deformations [Breton:2021mju, Ovgun:2019wej, Okyay:2021nnh]). In this context, the recent construction of new black hole solutions within both second- and first-order formulations of NLED sharpens the theoretical landscape by clarifying how distinct variational principles and constitutive structures map into physically inequivalent spacetime geometries, thereby enlarging the space of consistent strong-field models testable with gravitational and electromagnetic observations [Verbin:2024ewl], and complementing related strong-field extensions in modified gravity settings, including scalarized dyonic configurations [Brihaye:2021ich].
@@
These effects are not merely academic: magnetars, with surface magnetic fields of order $10^{14}$-$10^{15}\,$G, constitute natural laboratories where QED vacuum polarization is expected to leave measurable imprints on X-ray propagation and polarization [Duncan:1992hi, Kaspi:2017fwg]. The emergence of dedicated X-ray polarimetry missions has therefore transformed vacuum birefringence from a theoretical hallmark into a precision diagnostic, sharpening the need for theoretical predictions that correctly encode the relevant geometry of emission and detection [Weisskopf_2016, eXTP:2018anb]. In realistic magnetar observations, photons are emitted at finite radius (often from the stellar surface or magnetosphere) and are detected at large but finite distance, so the accumulated polarization-dependent bending and phase shift can differ systematically from idealized scattering calculations.
@@
A powerful way to formulate photon propagation in NLED is the effective (optical) metric approach. In the geometric-optics limit, the eikonal wave vector satisfies a Fresnel equation that generically factorizes into two polarization branches, implying that the two photon modes propagate along null geodesics of two distinct effective geometries [Boillat:1970gw, Plebanski:1959ff, Novello:1999pg, DeLorenci:2000yh]. In this picture, birefringence becomes a manifestly geometric phenomenon: it is encoded in the inequivalence of the corresponding optical manifolds, rather than introduced by hand through refractive indices. This viewpoint has been widely exploited to compute polarization-dependent deflection angles and related strong-field observables in various NLED models (including QED and Born-Infeld theory) under asymptotic assumptions [Bialynicka-Birula:1970nlh, Adler:1971wn, Born:1933pep, Born:1934gh]. Nonlinear electrodynamics (NED) generically modifies light propagation because electromagnetic perturbations propagate on an effective (field-dependent) geometry, leading to nontrivial causal and optical phenomena so strong, in fact, that closed, space-like photon paths can arise in suitable NED backgrounds [Novello:2001fv]. A systematic way to characterize these effects is provided by Fresnel (birefringence) analysis, which derives the dispersion relations and polarization-dependent light cones for general NED theories [Obukhov:2002xa]. In gravitational settings, NED also supports magnetically charged regular/“nonlinear magnetic” black holes whose dynamical viability can be tested via stability analyses [Breton:2005ye] and by studying their response to electromagnetic perturbations and late-time relaxation (ringdown) [Toshmatov:2018tyo, Toshmatov:2019gxg]. On the observational side, regular spacetimes such as Bardeen and Ayón–Beato–García exhibit distinctive circular-geodesic structure and disk phenomenology—impacting ISCO properties, profiled spectral lines from Keplerian disks/rings, and even producing “ghost” lensing images in no-horizon configurations [Stuchlik:2014qja, Schee:2016mjd, Schee:2015nua, Schee:2019gki].
@@
Strong field gravitational lensing provides a clean probe of the near horizon geometry, producing a sequence of relativistic images whose positions and magnifications encode the properties of the underlying compact object. The foundational strong lensing framework for Schwarzschild spacetime and its relativistic images was developed in Refs. [Virbhadra:1999nm, Virbhadra:2008ws]. Extending these ideas to nonlinear electrodynamics, Born–Infeld black holes were shown to yield distinctive lensing signatures relative to Maxwell electrovacuum [Eiroa:2005ag], and more generally light propagation can, in principle, discriminate the effective “charge” content of NED black holes through characteristic deviations in the deflection and optical observables [Toshmatov:2021fgm]. More recently, combining strong lensing observables with horizon scale imaging has emerged as a powerful avenue to constrain modified gravity models using both lensing data and black-hole shadow measurements [Kuang:2022ojj]. In particular, braneworld black holes exhibit characteristic modifications in their lensing and retrolensing observables compared to GR expectations, providing an astrophysical window into extra-dimensional corrections [Abdujabbarov:2017pfw]. Realistic propagation effects can further reshape these signatures: the presence of plasma alters photon trajectories and observable deflection, and this impact has been quantified for Schwarzschild black holes embedded in perfect-fluid dark matter environments [Atamurotov:2021hoq] as well as for Schwarzschild–MOG black holes, where both modified gravity and dispersive plasma effects jointly influence the weak-lensing regime [Atamurotov:2021qds]. Regular/quantum-corrected black holes provide a well-motivated arena where both shadow and weak lensing observables can deviate from the Schwarzschild/Kerr expectations, especially once realistic dispersive media are included. In asymptotically safe gravity, plasma-induced refraction can significantly reshape the apparent shadow size and distort weak-deflection predictions for regular black holes, strengthening the case for joint shadow–lensing tests in non-vacuum environments [Turakhonov:2025ojy]. Complementarily, charged black holes in regularized (4D) Einstein–Gauss–Bonnet gravity exhibit modified bending angles and image properties relative to GR, offering an additional probe of higher-curvature corrections through gravitational lensing [Kumar:2020sag]. Beyond modified gravity, nonlinear electrodynamics models can generate regular (and even horizonless ultracompact) geometries with distinct photon-sphere structure and shadow imprints, making shadow measurements a sharp discriminator among NED theories and compact-object classes [KumarWalia:2024yxn]. Recent work has sharpened the consistency requirements behind “regular black holes sourced by nonlinear electromagnetic fields,” emphasizing that singularity resolution must be confronted simultaneously with the NED field equations, global structure, and the effective (possibly birefringent) causal cones governing photon propagation [Bokulic:2025brf, Babaei-Aghbolagh:2025tim]. In strictly stationary settings and in broader analyses of NED-coupled spacetimes, nontrivial integrability and energy–momentum constraints can strongly restrict which NED models can genuinely support nonsingular geometries [Bokulic:2021xom, Bokulic:2022cyk], complementing earlier discussions on construction subtleties and consistency checks for “regular BH recipes” [Fan:2016hvf, Toshmatov:2018cks]. A central issue is stability: linear analyses for self-gravitating NED black holes and magnetic configurations reveal nontrivial mode conditions [Moreno:2002gg, Nomura:2020tpc], and recent results indicate that broad classes of NED nonsingular black holes can suffer instabilities, substantially constraining their viability as realistic endpoints of collapse [DeFelice:2024seu]. These theoretical constraints are further tightened by laboratory/astrophysical bounds on nonlinear corrections to Maxwell theory, including limits from light-by-light $($$\gamma\gamma$)) scattering and precision optical tests [NiauAkmansoy:2018ilv, Fouche:2016qqj]. Finally, the broader landscape of singularity resolution includes higher-derivative/gravity-sector mechanisms and alternative nonsingular compact objects (e.g., regular-center constructions and black-bounce-type solutions), offering complementary pathways beyond pure NED regularization [Cano:2020ezi, Maeda:2021jdc, Junior:2025sjr].
@@
In this work we develop such a finite-distance framework for vacuum birefringence in strong gravitational and electromagnetic fields. Our starting point is the Gauss-Bonnet method for light deflection, which expresses the bending angle as a global curvature integral over an appropriate domain in the optical manifold [Gibbons:2008rj]. We generalize this approach to the birefringent setting by constructing the two polarization-dependent optical geometries and by adopting the finite-distance prescription for geometric angles at the source and observer developed in the gravitational-lensing context [Ishihara:2016vdc, Ishihara:2016sfv, Rindler:2007zz]. This yields a clean, coordinate-invariant representation of the deflection for each polarization mode that remains valid when the endpoints are at finite radius.
@@
We apply the general formalism to two benchmark theories: the Euler-Heisenberg effective action of QED and the Born-Infeld model [Heisenberg:1936nmg, Schwinger:1951nm, Born:1933pep, Born:1934gh]. We derive closed expressions and controlled series expansions for the differential bending angle that explicitly retain the finite positions of the source and observer. Finally, specializing to magnetar-motivated configurations, we quantify the observational impact of finite-distance truncation by introducing a geometric suppression factor for surface (or near-surface) emission. We show that the birefringence predicted by asymptotic scattering models can be substantially reduced for realistic emission geometries, reaching order-unity corrections (up to $\sim 50\%$ in representative limb-emission configurations), and we emphasize that incorporating these effects is essential for faithful inference of magnetic-field strengths and NLED/QED parameters from X-ray polarimetry data [Weisskopf_2016, eXTP:2018anb].

### v2

@@ Abstract
We study polarization dependent photon propagation in static, spherically symmetric spacetimes permeated by strong magnetic fields, with the aim of quantifying how finite emission and detection radii modify vacuum birefringence signals. Working in the geometric optics limit of NLED, we formulate the two polarization modes as null geodesics of distinct effective (optical) metrics. We then develop a controlled weak-coupling expansion that cleanly separates the standard gravitational deflection from the birefringent contribution induced by the electromagnetic nonlinearity. Using a finite distance Gauss-Bonnet construction on the associated optical manifolds, we derive a general expression for the differential bending angle in which the source and observer are kept at arbitrary radii, thereby extending the usual scattering-at-infinity treatment. As benchmark applications, we specialize our results to the Euler-Heisenberg effective action of quantum electrodynamics (QED) and to Born-Infeld electrodynamics. We find that the observable birefringence is generically reduced by finite-distance truncation of the curvature flux, and we provide explicit correction series suitable for data analysis. For magnetar-motivated dipole-like falloff, the same geometric truncation can reduce the predicted polarization-dependent deflection at the order-tens-of-percent level for surface/near-surface emission; in a simple one-sided (outward-only) benchmark the suppression can approach $\sim 1/2$ for near-limb rays. Since realistic dipole magnetospheres are axisymmetric rather than spherically symmetric, we present this as an illustrative scaling estimate and leave a fully axisymmetric treatment to future work. Our results furnish a necessary finite-distance calibration for interpreting current and future X-ray polarimetry measurements and for placing unbiased constraints on strong-field QED and broader NLED parameters.
@@ I Introduction
The interaction of light with strong gravitational and electromagnetic fields provides a uniquely clean arena in which fundamental physics becomes observationally testable. In classical general relativity, light propagation is achromatic and polarization blind: null rays follow the spacetime light cone, and gravitational lensing is a purely geometric effect [35, 34]. Quantum field theory in external backgrounds, however, predicts that the vacuum itself acquires nontrivial optical properties in sufficiently intense electromagnetic fields. In particular, vacuum polarization in QED leads to the celebrated Euler-Heisenberg effective action and to birefringence and dispersion of photon modes in strong magnetic fields [43, 11, 2, 80]. Beyond QED, NLED also arises as an effective description in broader contexts, including string-inspired constructions, and admits a rich phenomenology when coupled to gravity [39, 90, 81, 51, 24, 28, 25].
@@
The advent of horizon-scale imaging by the Event Horizon Telescope (EHT), exemplified by the image of Sagittarius A*, has opened a new precision window for testing gravity and fundamental physics in the strong-field regime, where deviations from the Kerr paradigm can be constrained directly at the scale of the photon ring [92]. In this context, NLED provides a particularly well-motivated arena: magnetically charged black holes sourced by NLED can yield shadow and lensing signatures that are, in principle, accessible to EHT analyses and can be used to bound the underlying electromagnetic nonlinearity [3, 26]. More broadly, combining electromagnetic probes (shadow, lensing) with dynamical and particle-physics observables (quasinormal modes, greybody bounds, and propagation effects) offers a multi-messenger strategy to break degeneracies among model parameters [72, 56]. Recent studies of rotating Einstein-Euler-Heisenberg black holes further demonstrate how QED-inspired nonlinearities can imprint simultaneously on shadows and ringdown spectra, reinforcing the relevance of strong-field birefringent and dispersive effects for upcoming high-resolution observations [57]. NLED provides a well-motivated extension of Maxwell theory that captures high-field corrections arising both as effective quantum phenomena, such as the Euler-Heisenberg vacuum polarization effects underlying photon dispersion and photon splitting in strong backgrounds [11, 2], and as fundamental structures emerging in broader frameworks, including string-inspired constructions [39, 90]. When coupled to general relativity, NLED generically alters the near-horizon and strong-curvature regimes and has long been known to admit black-hole geometries whose properties can deviate markedly from the Reissner-Nordström solution [73, 31, 97, 76, 51]. A particularly influential development is that suitable NLED sources can yield regular (nonsingular) black holes, providing controlled arenas to explore the interplay between energy conditions, effective stress tensors, and astrophysical observables [7, 22, 33, 9, 37, 10, 58]. These geometries have renewed relevance because NLED-induced modifications can imprint themselves on wave and ray probes (e.g., quasinormal spectra, birefringence, and optical signatures such as weak deflection and shadow deformations [19, 71, 68]). In this context, the recent construction of new black hole solutions within both second- and first-order formulations of NLED sharpens the theoretical landscape by clarifying how distinct variational principles and constitutive structures map into physically inequivalent spacetime geometries, thereby enlarging the space of consistent strong-field models testable with gravitational and electromagnetic observations [93], and complementing related strong-field extensions in modified gravity settings, including scalarized dyonic configurations [21].
@@
These effects are not merely academic: magnetars, with surface magnetic fields of order $10^{14}$-$10^{15}\,$G, constitute natural laboratories where QED vacuum polarization is expected to leave measurable imprints on X-ray propagation and polarization [32, 47, 61, 84]. The emergence of dedicated X-ray polarimetry missions has therefore transformed vacuum birefringence from a theoretical hallmark into a precision diagnostic, sharpening the need for theoretical predictions that correctly encode the relevant geometry of emission and detection [96, 98, 85, 55]. In realistic magnetar observations, photons are emitted at finite radius (often from the stellar surface or magnetosphere) and are detected at large but finite distance, so the accumulated polarization-dependent bending and phase shift can differ systematically from idealized scattering calculations.
@@
A powerful way to formulate photon propagation in NLED is the effective (optical) metric approach. In the geometric-optics limit, the eikonal wave vector satisfies a Fresnel equation that generically factorizes into two polarization branches, implying that the two photon modes propagate along null geodesics of two distinct effective geometries [13, 74, 65, 30]. In this picture, birefringence becomes a manifestly geometric phenomenon: it is encoded in the inequivalence of the corresponding optical manifolds, rather than introduced by hand through refractive indices. This viewpoint has been widely exploited to compute polarization-dependent deflection angles and related strong-field observables in birefringent NLED models (notably QED-type effective actions) under asymptotic assumptions [11, 2, 17, 18]. Exceptional NLEDs also exist: in particular, Born-Infeld electrodynamics is non-birefringent–the two polarization modes propagate on the same effective light cone (up to conformal rescaling) [13, 67]. NLED generically modifies light propagation because electromagnetic perturbations propagate on an effective (field-dependent) geometry, leading to nontrivial causal and optical phenomena so strong, in fact, that closed, space-like photon paths can arise in suitable NLED backgrounds [66]. A systematic way to characterize these effects is provided by Fresnel (birefringence) analysis, which derives the dispersion relations and polarization-dependent light cones for general NLED theories [67]. In gravitational settings, NLED also supports magnetically charged regular/“nonlinear magnetic” black holes whose dynamical viability can be tested via stability analyses [20] and by studying their response to electromagnetic perturbations and late-time relaxation (ringdown) [89, 87]. On the observational side, regular spacetimes such as Bardeen and Ayón–Beato–García exhibit distinctive circular-geodesic structure and disk phenomenology—impacting ISCO properties, profiled spectral lines from Keplerian disks/rings, and even producing “ghost” lensing images in no-horizon configurations [82, 77, 78, 79].
@@
Strong field gravitational lensing provides a clean probe of the near horizon geometry, producing a sequence of relativistic images whose positions and magnifications encode the properties of the underlying compact object. The foundational strong lensing framework for Schwarzschild spacetime and its relativistic images was developed in Refs. [94, 95]. Extending these ideas to NLED, Born–Infeld black holes were shown to yield distinctive lensing signatures relative to Maxwell electrovacuum [36], and more generally light propagation can, in principle, discriminate the effective “charge” content of NLED black holes through characteristic deviations in the deflection and optical observables [86]. More recently, combining strong lensing observables with horizon scale imaging has emerged as a powerful avenue to constrain modified gravity models using both lensing data and black-hole shadow measurements [52]. In particular, braneworld black holes exhibit characteristic modifications in their lensing and retrolensing observables compared to GR expectations, providing an astrophysical window into extra-dimensional corrections [1]. Realistic propagation effects can further reshape these signatures: the presence of plasma alters photon trajectories and observable deflection, and this impact has been quantified for Schwarzschild black holes embedded in perfect-fluid dark matter environments [5] as well as for Schwarzschild–MOG black holes, where both modified gravity and dispersive plasma effects jointly influence the weak-lensing regime [6]. Regular/quantum-corrected black holes provide a well-motivated arena where both shadow and weak lensing observables can deviate from the Schwarzschild/Kerr expectations, especially once realistic dispersive media are included. In asymptotically safe gravity, plasma-induced refraction can significantly reshape the apparent shadow size and distort weak-deflection predictions for regular black holes, strengthening the case for joint shadow–lensing tests in non-vacuum environments [91]. Complementarily, charged black holes in regularized (4D) Einstein–Gauss–Bonnet gravity exhibit modified bending angles and image properties relative to GR, offering an additional probe of higher-curvature corrections through gravitational lensing [53]. Beyond modified gravity, NLED models can generate regular (and even horizonless ultracompact) geometries with distinct photon-sphere structure and shadow imprints, making shadow measurements a sharp discriminator among NLED theories and compact-object classes [54]. Recent work has sharpened the consistency requirements behind “regular black holes sourced by nonlinear electromagnetic fields,” emphasizing that singularity resolution must be confronted simultaneously with the NLED field equations, global structure, and the effective (possibly birefringent) causal cones governing photon propagation [15, 8]. In strictly stationary settings and in broader analyses of NLED-coupled spacetimes, nontrivial integrability and energy–momentum constraints can strongly restrict which NLED models can genuinely support nonsingular geometries [14, 16], complementing earlier discussions on construction subtleties and consistency checks for “regular BH recipes” [37, 88]. A central issue is stability: linear analyses for self-gravitating NLED black holes and magnetic configurations reveal nontrivial mode conditions [62, 64], and recent results indicate that broad classes of NLED nonsingular black holes can suffer instabilities, substantially constraining their viability as realistic endpoints of collapse [29]. These theoretical constraints are further tightened by laboratory/astrophysical bounds on nonlinear corrections to Maxwell theory, including limits from light-by-light $($$\gamma\gamma$)) scattering and precision optical tests [63, 38]. Finally, the broader landscape of singularity resolution includes higher-derivative/gravity-sector mechanisms and alternative nonsingular compact objects (e.g., regular-center constructions and black-bounce-type solutions), offering complementary pathways beyond pure NLED regularization [23, 60, 46].
@@
In this work we develop such a finite-distance framework for vacuum birefringence in strong gravitational and electromagnetic fields. Our starting point is the Gauss-Bonnet method for light deflection, which expresses the bending angle as a global curvature integral over an appropriate domain in the optical manifold [41]. We generalize this approach to the birefringent setting by constructing the two polarization-dependent optical geometries and by adopting the finite-distance prescription for geometric angles at the source and observer developed in the gravitational-lensing context [45, 44, 75]. This yields a clean, coordinate-invariant representation of the deflection for each polarization mode that remains valid when the endpoints are at finite radius.
@@
We apply the general formalism to two benchmark theories: the Euler-Heisenberg effective action of QED and Born-Infeld (BI) electrodynamics [43, 80, 17, 18], used here as a non-birefringent consistency check. We derive closed expressions and controlled series expansions for the differential bending angle that explicitly retain the finite positions of the source and observer. In particular, with our definition of vacuum birefringence as the differential deflection $\Delta\alpha\equiv\alpha_{+}-\alpha_{-}$, Born-Infeld theory yields identically $\Delta\alpha_{\rm BI}=0$ (the two polarization modes share a single characteristic cone) [13, 67]. Finally, specializing to magnetar-motivated configurations, we quantify the observational impact of finite-distance truncation by introducing a geometric suppression factor for surface (or near-surface) emission. We show that the birefringence predicted by asymptotic scattering models can be substantially reduced for realistic emission geometries, reaching order-unity corrections (up to $\sim 50\%$ in representative limb-emission configurations), and we emphasize that incorporating these effects is essential for faithful inference of magnetic-field strengths and birefringent NLED/QED parameters from X-ray polarimetry data [96, 98].